# Revealing the building blocks of tree balance: fundamental units of the Sackin and Colless Indices

### v1

@@ 3.1.2 Minimizing $\Delta_{CS}$ and maximizing $N_{b}$ as well as $\Delta_{SC}$
We now assume that uniqueness is not guaranteed, i.e., that there are powers of 2 for which the fully balanced tree is not the unique minimizer of $\Delta_{CS}$. Let $n=2^{k_{n}}$ be minimal with this property, and let $T=(T_{a},T_{b})$ be a tree with $n$ leaves which is not fully balanced but minimizes $\Delta_{CS}$ (and thus maximizes $N_{b}$ and $\Delta_{SC}$).
@@
We continue with a full characterization of all $\Delta_{CS}$ minimal (and thus $N_{b}$ and $\Delta_{SC}$ maximal) trees. Note that while we know from Theorem 3 that all Colless minima also minimize $\Delta_{CS}$, and by Proposition 4 that the same is true for the echelon tree, these trees generally form only a subset of $\Delta_{CS}$ minimal trees, cf. Figure 3.
Figure 3: A depiction of all $\Delta_{CS}$ minimal (and $\Delta_{SC}$ as well as $N_{b}$ maximal) trees for $n=9$: The set consists of $T_{9}^{gfb}$, $T_{9}^{be}$ and $T=(T_{a},T_{b})$. Note that at the root, $T$ has a Colless partition as $n_{a}=n_{a}^{gfb}=5$ and $n_{b}=n_{b}^{gfb}=4$. However, $T$ is not Colless minimal as it employs a non-Colless minimal subtree $T_{a}=T_{5}^{be}$. As this subtree is $\Delta_{CS}$ minimal, too, so is $T$. Note that $T$ can be regarded as a “combination” of the echelon minimality property and the Colless minimality property. Theorem 5 shows that echelon trees, Colless minima and combinations thereof are the only $\Delta_{CS}$ minima and $N_{b}$ as well as $\Delta_{SC}$ maxima.
However, it turns out that these trees are the basic building blocks of which all $\Delta_{CS}$ minimal and both $N_{b}$ and $\Delta_{SC}$ maximal trees consist. Our characterization of $\Delta_{CS}$ minimal and thus $N_{b}$ and $\Delta_{SC}$ maximal trees is based on pairs $(n_{a},n_{b})$ which are either Colless partitions as defined in Section 2.2.4 or echelon partitions as defined in Section 2.1.
@@
We show both directions of the “if and only if”-statement separately.
@@
    We first want to show that every $\Delta_{CS}$ minimal tree has the property stated by the theorem. Seeking a contradiction, we assume this is not the case, i.e., assume there are trees that minimize $\Delta_{CS}$ but have an inner node whose induced subtree’s standard decomposition does not lead to a Colless or echelon partition. We consider a smallest $n$ for which such a tree $T=(T_{a},T_{b})$ exists and note that $T_{a}$ and $T_{b}$ must fulfill the conditions of the theorem as they are also $\Delta_{CS}$ minimal by Corollary 1 and as they have fewer than $n$ leaves and $n$ was chosen to be minimal. Let $T_{i}$ have $n_{i}$ many leaves for $i=\{a,b\}$ with $n_{a}\geq n_{b}$. We know that $(n_{a},n_{b})$ is neither a Colless nor an echelon partition and we will show that this leads to a contradiction. In the following, we again define $k_{n}=\lceil\log_{2}n\rceil$.
    We consider three cases now based on the difference $n_{a}-n_{b}$.
@@
      Consider the case $n_{a}-n_{b}>2^{k_{n}}-n=2^{k_{n}-1}-(n-2^{k_{n}-1})=n_{a}^{be}-n_{b}^{be}$, where $(n_{a}^{be},n_{b}^{be})=(2^{k_{n}-1},n-2^{k_{n}-1})$ is the echelon partition. As $n_{a}+n_{b}=n_{a}^{be}+n_{b}^{be}$, this must imply $n_{a}>n_{a}^{be}=2^{k_{n}-1}$. This, however, shows, by Lemma 4, that $T$ is not $\Delta_{CS}$ minimal (and not $N_{b}$ or $\Delta_{SC}$ maximal). This is a contradiction to our choice of $T$, implying this scenario cannot happen.
@@
      Consider the case $n_{a}-n_{b}\leq n_{a}^{gfb}-n_{b}^{gfb}$, where $(n_{a}^{gfb},n_{b}^{gfb})$ is the partition of $n$ as induced by the greedy from the bottom tree $T_{n}^{gfb}$, cf. Section 2.2.1. We know that $(n_{a},n_{b})$ is a Sackin partition, cf. Section 2.2.3. Since we assume $(n_{a},n_{b})$ is not a Colless partition, it must be true that $(n_{a},n_{b})$ is a partition that has a corresponding Sackin minimal tree, but not a Colless minimal tree. We now construct a tree $T^{\prime}=(T_{a}^{\prime},T_{b}^{\prime})$ as follows: $T_{a}^{\prime}$ has $n_{a}^{\prime}=n_{a}$ leaves and $T_{a}^{\prime}$ has $n_{b}^{\prime}=n_{b}$ leaves, and $T_{a}^{\prime}$ and $T_{b}^{\prime}$ are both Colless minima (and thus also Sackin minima). This way, we ensure that $T^{\prime}$ is a Sackin minimum and that $T_{a}^{\prime}$ and $T_{b}^{\prime}$ are also $\Delta_{CS}$ minimal by Theorem 3.
@@
      Using Proposition 1, we now argue that $\Delta_{CS}(T)=\Delta_{CS}(T^{\prime})$:
@@
      $=\Delta_{CS}(T_{a}^{\prime})+\Delta_{CS}(T_{b}^{\prime})-2n_{b}^{\prime}$
      $=\Delta_{CS}(T^{\prime}).$
      Next, let $T^{\prime\prime}=(T_{a}^{\prime\prime},T_{b}^{\prime\prime})$ with $n_{a}^{\prime\prime}$, $n_{b}^{\prime\prime}$ leaves and $n_{a}^{\prime\prime}\geq n_{b}^{\prime\prime}$ be a Colless minimal tree on $n$ leaves. As by Theorem 3, $T^{\prime\prime}$ is also $\Delta_{CS}$ minimal (and $N_{b}$ as well as $\Delta_{SC}$ maximal), we know $\Delta_{CS}(T)=\Delta_{CS}(T^{\prime})=\Delta_{CS}(T^{\prime\prime})$. We will show that the latter equality is a contradiction.
@@
      where the inequality stems from the fact that $T^{\prime\prime}$ is a Colless minimum, whereas $T^{\prime}$ cannot be one (as $(n_{a}^{\prime},n_{b}^{\prime})=(n_{a},n_{b})$ is not a Colless partition of $n$). The first equality sign after the inequality is due to the fact that $T^{\prime}$ is a Sackin minimum and $T^{\prime\prime}$ – as a Colless minimum – is guaranteed to be a Sackin minimum as well. Thus, we must have $S(T^{\prime})=S(T^{\prime\prime})$. However, this shows that the equality $\Delta_{CS}(T^{\prime})=\Delta_{CS}(T^{\prime\prime})$ as derived above is a contradiction, implying this scenario cannot happen.
    3. (c)
      Consider the case $n_{a}^{gfb}-n_{b}^{gfb}<n_{a}-n_{b}<n_{a}^{be}-n_{b}^{be}=2^{k_{n}-1}-n_{b}^{be}$. Note that by $n=n_{a}^{gfb}+n_{b}^{gfb}=n_{a}+n_{b}=2^{k_{n}-1}+n_{b}^{be}$, the assumed inequalities imply in particular that $n_{a}^{gfb}<n_{a}<2^{k_{n}-1}$. We will use these inequalities subsequently as we consider two subcases.
      1. i.
        If $n\geq 3\cdot 2^{k_{n}-2}$, we know that $n_{a}^{gfb}=2^{k_{n}-1}$ and $n_{b}^{gfb}=n-2^{k_{n}-1}$, cf. Section 2.2.1. Thus, we have $n_{a}^{gfb}-n_{b}^{gfb}=2^{k_{n}}-n$.
        Now consider $n_{a}-n_{b}=n_{a}-(n-n_{a})=2n_{a}-n<2^{k_{n}}-n=n_{a}^{gfb}-n_{b}^{gfb}$. Here, the inequality sign stems from the fact that we have already shown $n_{a}<2^{k_{n}-1}$. However, $n_{a}-n_{b}<n_{a}^{gfb}-n_{b}^{gfb}$ is a contradiction to our assumption in the current case.
      2. ii.
        If $n<3\cdot 2^{k_{n}-2}$, we know $n_{b}^{gfb}=2^{k_{n}-2}$, cf. Section 2.2.1, and, by $n_{a}-n_{b}>n_{a}^{gfb}-n_{b}^{gfb}$ as stated above, we have $2^{k_{n}-1}>n_{a}>n_{a}^{gfb}\geq n_{b}^{gfb}=2^{k_{n}-2}>n_{b}$. This shows that $T_{n_{a}}^{be}$, the echelon tree with $n_{a}$ leaves, must contain $T_{k_{n}-2}^{fb}$ as a maximum pending subtree (as $2^{k_{n}-2}$ is the largest power of two smaller than $n_{a}$). We now construct tree $\widetilde{T}=(T_{n_{a}}^{be},T_{b})$ as depicted in Figure 4. This tree has the same maximum pending subtree $T_{b}$ as $T$, but $T_{a}$ has been replaced by $T_{n_{a}}^{be}$. However, since $T$ is $\Delta_{CS}$ minimal, $T_{a}$ is $\Delta_{CS}$ minimal by Corollary 1. Furthermore, $T_{n_{a}}^{be}$ is also $\Delta_{CS}$ minimal by Proposition 4 as it is an echelon tree. Therefore, we must have $\Delta_{CS}(T_{a})=\Delta_{CS}(T_{n_{a}}^{be})$. This leads to:
@@
        Figure 4: Trees $\widetilde{T}$ and $T^{\prime}$ as needed in the proof of Theorem 5. Note that both trees employ the same pending subtrees and the only vertices in which they differ are highlighted with diamonds.
        We now consider tree $T^{\prime}$ from Figure 4. $T^{\prime}$ is like $\widetilde{T}$, but $T_{b}$ and $T_{k_{n}-2}^{fb}$ are interchanged. Note that $\widetilde{T}$ and $T^{\prime}$ differ only in the two vertices marked with diamonds. This immediately leads to:
        $C(T^{\prime})$ $=C(\widetilde{T})-|(n_{a}-2^{k_{n}-2})-2^{k_{n}-2}|-|(n_{a}-2^{k_{n}-2}+2^{k_{n}-2})-n_{b}|$
        $\ \ +|(n_{a}-2^{k_{n}-2})-n_{b}|+|(n_{a}-2^{k_{n}-2}+n_{b})-2^{k_{n}-2}|$
        $=C(\widetilde{T})-|n_{a}-2^{k_{n}-1}|-|n_{a}-n_{b}|+|n_{a}-n_{b}-2^{k_{n}-2}|+|n_{a}+n_{b}-2^{k_{n}-1}|$
        $=C(\widetilde{T})-(2^{k_{n}-1}-n_{a})-n_{a}+n_{b}+|n_{a}-n_{b}-2^{k_{n}-2}|+n_{a}+n_{b}-2^{k_{n}-1}$
        $=C(\widetilde{T})-2^{k_{n}}+n_{a}+2n_{b}+|n_{a}-n_{b}-2^{k_{n}-2}|$
        $S(T^{\prime})$ $=S(\widetilde{T})-(n_{a}-2^{k_{n}-2}+2^{k_{n}-2})+(n_{a}-2^{k_{n}-2}+n_{b})$
        $=S(\widetilde{T})-2^{k_{n}-2}+n_{b},$
        which directly shows:
        $\Delta_{CS}(T^{\prime})$ $=C(T^{\prime})-S(T^{\prime})$
        $=(C(\widetilde{T})-2^{k_{n}}+n_{a}+2n_{b}+|n_{a}-n_{b}-2^{k_{n}-2}|)-(S(\widetilde{T})-2^{k_{n}-2}+n_{b})$
        $=C(\widetilde{T})-S(\widetilde{T})+n_{a}+n_{b}-3\cdot 2^{k_{n}-2}+|n_{a}-n_{b}-2^{k_{n}-2}|$
        $=\Delta_{CS}(\widetilde{T})+n_{a}+n_{b}-3\cdot 2^{k_{n}-2}+|n_{a}-n_{b}-2^{k_{n}-2}|.$ (5)
        As we do not know if the echelon subtree with $n_{a}-2^{k_{n}-2}$ many leaves has at least as many leaves as $T_{b}$ with $n_{b}$ leaves, we have to consider both cases for $|n_{a}-n_{b}-2^{k_{n}-2}|$. If the echelon tree is as least as large as $T_{b}$, we get:
[…]
        $=\Delta_{CS}(\widetilde{T})+2n_{a}-2^{k_{n}}$
[…]
        where the latter inequality stems from the fact that $n_{a}<2^{k_{n}-1}$.
        If, however, the echelon subtree is smaller than $T_{b}$, we get:
[…]
        $=\Delta_{CS}(\widetilde{T})+2n_{b}-2^{k_{n}-1}$
[…]
        where the latter inequality stems from the fact that $n_{b}<2^{k_{n}-2}$.
[…]
    As all cases lead to contradictions, the assumption was wrong and such a tree $T$ cannot exist. This completes the first part of the proof for $\Delta_{CS}$. The respective statement for $\Delta_{SC}$ then follows from $\Delta_{CS}=-\Delta_{SC}$, and the one for $N_{b}$ follows from Theorem 3.
  2. 2.
    Next, we want to show that every tree that at each of its inner vertices has a pending subtree inducing a Colless or echelon partition is indeed $\Delta_{CS}$ minimal. Suppose this is not the case for all such trees and consider $T=(T_{a},T_{b})$ to be a minimal counterexample, i.e., a counterexample with the minimal number $n$ of leaves. Let $n_{a}$ denote the number of leaves of $T_{a}$ and $n_{b}$ the number of leaves of $T_{b}$ with $n_{a}\geq n_{b}$. Then, as $T$ has the property that all its inner vertices induce Colless or echelon partitions, so do $T_{a}$ and $T_{b}$. However, as $n_{a},n_{b}<n$ and by the minimality choice of $n$, we know that $T_{a}$ and $T_{b}$ are in fact $\Delta_{CS}$ minimal.
    Now we choose $T_{a}^{\prime}$ and $T_{b}^{\prime}$ as follows: If $(n_{a},n_{b})$ is a Colless partition, we choose $T_{a}^{\prime}$ and $T^{\prime}_{b}$ to be Colless minimal trees with $n_{a}$ and $n_{b}$ leaves, respectively. If, on the other hand, $(n_{a},n_{b})$ is the echelon partition, we choose $T_{a}^{\prime}$ and $T^{\prime}_{b}$ to be echelon trees with $n_{a}$ and $n_{b}$ leaves, respectively. We now consider $T^{\prime}=(T_{a}^{\prime},T_{b}^{\prime})$. As $T_{a}^{\prime}$ and $T^{\prime}_{b}$ are either Colless minima or echelon trees, they are $\Delta_{CS}$ minimal (and $N_{b}$ as well as $\Delta_{SC}$ maximal) by Theorem 3 or Proposition 4, respectively. Hence, by the minimality of $T_{a}$ and $T_{b}$, we must have $\Delta_{CS}(T_{a})=\Delta_{CS}(T_{a}^{\prime})$ and $\Delta_{CS}(T_{b})=\Delta_{CS}(T_{b}^{\prime})$. This immediately leads to:
[…]
    $=\Delta_{CS}(T_{a})+\Delta_{CS}(T_{b})-2n_{b}$
    $=\Delta_{CS}(T).$
    Thus, we have $\Delta_{CS}(T)=\Delta_{CS}(T^{\prime})$. However, $T^{\prime}$ is either a Colless minimum or an echelon tree, as its maximum pending subtrees have the respective properties and as its partition of $n$ at the root was consistent with either being Colless minimal or echelon, respectively. Thus, we know – again by Theorem 3 or Proposition 4, respectively – that $T^{\prime}$ is $\Delta_{CS}$ minimal. This contradicts the assumption that $T$ is not $\Delta_{CS}$ minimal and thus completes the proof. The respective statement for $N_{b}$ follows then again from Lemma 2, and the one for $\Delta_{SC}$ from $\Delta_{CS}=-\Delta_{SC}$.
We conclude this section with the following corollary, which gives a recursive formula to count the number of $\Delta_{CS}$ minima and $N_{b}$ as well as $\Delta_{SC}$ maxima.
@@
The proof is a direct consequence of Theorem 5 as all minimal trees are composed of two maximal pending subtrees which are also minimal, and these must have sizes that induce Colless or echelon partitions. In the even case, we need to avoid counting a combination of trees twice, which is why the simple sum is only considering the case $n_{a}>n_{b}$, whereas the case $n_{a}=n_{b}$ is addressed by the binomial coefficient for trees $T=(T_{a},T_{b})$ with $n_{a}=n_{b}=\frac{n}{2}$ and $T_{a}\neq T_{b}$, and by the term $a\left(\frac{n}{2}\right)$ for trees $T=(T_{a},T_{b})$ with $T_{a}=T_{b}$. This completes the proof. ∎
We conclude this section by noting that the sequence $(a_{n})_{n\geq 1}$ with $a_{n}=a(n)$, which starts with the values $\{1,1,1,1,2,2,1,1,3,6,6,4,4,3,1,1,4,13,25,26\}$ was not contained in the Online Encyclopedia of Integer Sequences OEIS [24], which indicates that it might not have appeared in other contexts before. It has been submitted to the OEIS in the course of this manuscript.
3.1.3 Other properties of $\Delta_{CS}$, $\Delta_{SC}$ and $N_{b}$

### v2

@@ 3.1.2 Minimizing $\Delta_{CS}$ and maximizing $N_{b}$ as well as $\Delta_{SC}$
We now assume that uniqueness is not guaranteed, i.e., that there are powers of 2 for which the fully balanced tree is not the unique minimizer of $\Delta_{CS}$. Let $n=2^{k_{n}}$ be minimal with this property, and let $T=(T_{a},T_{b})$ be a tree with $n$ leaves which is not fully balanced but minimizes $\Delta_{CS}$.
@@
Before we can characterize all $\Delta_{CS}$ minimal trees, we need one more lemma.
Lemma 5.
Let $T=(T_{a},T_{b})$ be a rooted binary tree with $n=n_{a}+n_{b}$ leaves with $n_{a}\geq n_{b}$ being the numbers of leaves in $T_{a}$ and $T_{b}$, respectively. Let $k_{n}=\lceil\log_{2}(n)\rceil$ and let $(n_{a}^{gfb},n_{b}^{gfb})$ and $(n_{a}^{be}=2^{k_{n}-1},n_{b}^{be})$ denote the gfb and echelon partitions, respectively. Then, if $n_{a}^{gfb}<n_{a}<n_{a}^{be}=2^{k_{n}-1}$, $T$ is not $\Delta_{CS}$ minimal.
Proof.
Seeking a contradiction, we assume there is a tree $T=(T_{a},T_{b})$ with $n_{a}^{gfb}<n_{a}<n_{a}^{be}=2^{k_{n}-1}$ that is $\Delta_{CS}$ minimal. First note that if $n\geq 3\cdot 2^{k_{n}-2}$, we know by Section 2 that $n_{a}^{gfb}=2^{k_{n}-1}=n_{a}^{be}$, which shows that then $n_{a}^{gfb}<n_{a}<n_{a}^{be}$ cannot happen. So we must have $n<3\cdot 2^{k_{n}-2}$. Then, we have $n_{b}^{gfb}=2^{k_{n}-2}$ and thus $2^{k_{n}-1}>n_{a}>n_{a}^{gfb}\geq n_{b}^{gfb}=2^{k_{n}-2}>n_{b}$. This shows that $T_{n_{a}}^{be}$, the echelon tree with $n_{a}$ leaves, must contain $T_{k_{n}-2}^{fb}$ as a maximum pending subtree. We now construct tree $\widetilde{T}=(T_{n_{a}}^{be},T_{b})$ as depicted in Figure 3. This tree has the same maximum pending subtree $T_{b}$ as $T$, but $T_{a}$ has been replaced by $T_{n_{a}}^{be}$. However, since $T$ is $\Delta_{CS}$ minimal, $T_{a}$ is $\Delta_{CS}$ minimal by Corollary 1. Furthermore, $T_{n_{a}}^{be}$ is also $\Delta_{CS}$ minimal by Proposition 4 as it is an echelon tree. Therefore, we must have $\Delta_{CS}(T_{a})=\Delta_{CS}(T_{n_{a}}^{be})$. This leads to:
[…]
Figure 3: Trees $\widetilde{T}$ and $T^{\prime}$ as needed in the proof of Lemma 5. Note that both trees employ the same pending subtrees and the only vertices in which they differ are highlighted with diamonds.
We now consider tree $T^{\prime}$ from Figure 3. $T^{\prime}$ is like $\widetilde{T}$, but $T_{b}$ and $T_{k_{n}-2}^{fb}$ are interchanged. Note that $\widetilde{T}$ and $T^{\prime}$ differ only in the two vertices marked with diamonds. This immediately leads to (calculations omitted as they are analogous to those of Lemma 4):
$C(T^{\prime})$ $=C(\widetilde{T})-2^{k_{n}}+n_{a}+2n_{b}+|n_{a}-n_{b}-2^{k_{n}-2}|$
$S(T^{\prime})$ $=S(\widetilde{T})-2^{k_{n}-2}+n_{b},$
which directly shows:
$\Delta_{CS}(T^{\prime})$ $=C(T^{\prime})-S(T^{\prime})$
$=\Delta_{CS}(\widetilde{T})+n_{a}+n_{b}-3\cdot 2^{k_{n}-2}+|n_{a}-n_{b}-2^{k_{n}-2}|$
$=\Delta_{CS}(\widetilde{T})+n_{a}+n_{b}-3\cdot 2^{k_{n}-2}+\max\{n_{a}-n_{b}-2^{k_{n}-2},-n_{a}+n_{b}+2^{k_{n}-2}\}$
[…]
We continue with a full characterization of all $\Delta_{CS}$ minimal trees. Note that while we know from Theorem 3 that all Colless minima also minimize $\Delta_{CS}$, and by Proposition 4 that the same is true for the echelon tree, these trees generally form only a subset of $\Delta_{CS}$ minimal trees, cf. Figure 4.
Figure 4: A depiction of all $\Delta_{CS}$ minimal trees for $n=9$: The set consists of $T_{9}^{gfb}$, $T_{9}^{be}$ and $T=(T_{a},T_{b})$. Note that at the root, $T$ has a Colless partition as $n_{a}=n_{a}^{gfb}=5$ and $n_{b}=n_{b}^{gfb}=4$. However, $T$ is not Colless minimal as it employs a non-Colless minimal subtree $T_{a}=T_{5}^{be}$. As this subtree is $\Delta_{CS}$ minimal, too, so is $T$. Note that $T$ can be regarded as a “combination” of the echelon minimality property and the Colless minimality property. Theorem 6 shows that echelon trees, Colless minima and combinations thereof are the only $\Delta_{CS}$ minima.
However, it turns out that these trees are the basic building blocks of which all $\Delta_{CS}$ minimal trees consist. Our characterization of $\Delta_{CS}$ minimal trees is based on pairs $(n_{a},n_{b})$ which are either Colless partitions as defined in Section 2.2.4 or echelon partitions as defined in Section 2.1.
@@
We show the directions of the “if and only if”-statement separately.
@@
    We first want to show that every $\Delta_{CS}$ minimal tree has the property stated by the theorem. Seeking a contradiction, we consider a smallest $n$ for which such a $T=(T_{a},T_{b})$ exists which minimizes $\Delta_{CS}$ but which has an inner node that is neither the echelon nor a Colless partition. By the minimality of $T$, this node must be the root, which shows that $(n_{a},n_{b})$ is neither a Colless nor an echelon partition (where $n_{a}$, $n_{b}$ denote the numbers of leaves of $T_{a}$, $T_{b}$). Note that $T_{a}$ and $T_{b}$ must fulfill the conditions of the theorem as they are also $\Delta_{CS}$ minimal by Corollary 1 and as they have fewer than $n$ leaves. In the following, we again define $k_{n}=\lceil\log_{2}n\rceil$.
    We consider three cases now based on $n_{a}$.
@@
      Consider the cases $n_{a}>2^{k_{n}-1}=n_{a}^{be}$ or $n_{a}^{gfb}<n_{a}<n_{a}^{be}=2^{k_{n}-1}$. Then, by Lemma 4 or Lemma 5, respectively, $T$ is not $\Delta_{CS}$ minimal, a contradiction.
@@
      Note that the only remaining case is $n_{a}<n_{a}^{gfb}$. This is due to the fact that we have excluded all values of $n_{a}$ larger than $n_{a}^{gfb}$ except for $n_{a}^{be}$, but $n_{a}^{be}$ and $n_{a}^{gfb}$ are additionally excluded by the assumption that $(n_{a},n_{b})$ is neither a Colless nor an echelon partition. Thus, the only remaining case we need to analyze is the case $n_{a}<n_{a}^{gfb}$.
      In this case, we know that $(n_{a},n_{b})$ is a Sackin partition, cf. Section 2.2.3. Since we assume $(n_{a},n_{b})$ is not a Colless partition, it must be true that $(n_{a},n_{b})$ is a partition that has a corresponding Sackin minimal tree, but not a Colless minimal tree. We now construct a tree $T^{\prime}=(T_{a}^{\prime},T_{b}^{\prime})$ with $n$ leaves as follows: $T_{a}^{\prime}$ has $n_{a}^{\prime}=n_{a}$ leaves and $T_{b}^{\prime}$ has $n_{b}^{\prime}=n_{b}$ leaves, and $T_{a}^{\prime}$ and $T_{b}^{\prime}$ are both Colless minima (and thus also Sackin minima). This way, we ensure that $T^{\prime}$ is a Sackin minimum and that $T_{a}^{\prime}$ and $T_{b}^{\prime}$ are also $\Delta_{CS}$ minimal by Theorem 3.
@@
      Using Proposition 1, we can conclude that $\Delta_{CS}(T)=\Delta_{CS}(T^{\prime})$:
@@
      Next, let $T^{\prime\prime}=(T_{a}^{\prime\prime},T_{b}^{\prime\prime})$ with $n=n_{a}^{\prime\prime}+n_{b}^{\prime\prime}$ leaves and $n_{a}^{\prime\prime}\geq n_{b}^{\prime\prime}$ be a Colless minimal tree. As by Theorem 3, $T^{\prime\prime}$ is also $\Delta_{CS}$ minimal, we know $\Delta_{CS}(T)=\Delta_{CS}(T^{\prime})=\Delta_{CS}(T^{\prime\prime})$. We now show that the latter equality is a contradiction:
@@
      where the inequality stems from the fact that $T^{\prime\prime}$ is a Colless minimum, whereas $T^{\prime}$ cannot be one (as $(n_{a}^{\prime},n_{b}^{\prime})=(n_{a},n_{b})$ is not a Colless partition of $n$). The first equality sign after the inequality is due to the fact that $T^{\prime}$ and $T^{\prime\prime}$ are both Sackin minimal trees, implying $S(T^{\prime})=S(T^{\prime\prime})$. However, this shows that the equality $\Delta_{CS}(T)=\Delta_{CS}(T^{\prime\prime})$ as derived above is a contradiction.
    As all cases lead to contradictions, the assumption was wrong and such a tree $T$ cannot exist. This completes the first part of the proof.
  2. 2.
    Next, we want to show that every tree that at each of its inner vertices has a pending subtree inducing a Colless or echelon partition is indeed $\Delta_{CS}$ minimal. Suppose this is not the case for all such trees and consider $T=(T_{a},T_{b})$ to be a counterexample with the minimal number $n$ of leaves. Let $n_{a}$ denote the number of leaves of $T_{a}$ and $n_{b}$ the number of leaves of $T_{b}$ with $n_{a}\geq n_{b}$. Then, as $T$ has the property that all its inner vertices induce Colless or echelon partitions, so do $T_{a}$ and $T_{b}$. However, as $n_{a},n_{b}<n$ and by the minimality choice of $n$, we know that $T_{a}$ and $T_{b}$ are in fact $\Delta_{CS}$ minimal. Now we choose $T_{a}^{\prime}$ and $T_{b}^{\prime}$ as follows: If $(n_{a},n_{b})$ is a Colless partition, we choose $T_{a}^{\prime}$ and $T^{\prime}_{b}$ to be Colless minimal trees with $n_{a}$ and $n_{b}$ leaves, respectively. If, on the other hand, $(n_{a},n_{b})$ is the echelon partition, we choose $T_{a}^{\prime}$ and $T^{\prime}_{b}$ to be echelon trees with $n_{a}$ and $n_{b}$ leaves, respectively. We now consider $T^{\prime}=(T_{a}^{\prime},T_{b}^{\prime})$. As $T_{a}^{\prime}$ and $T^{\prime}_{b}$ are either Colless minima or echelon trees, they are $\Delta_{CS}$ minimal by Theorem 3 or Proposition 4, respectively. Hence, by the minimality of $T_{a}$ and $T_{b}$, we must have $\Delta_{CS}(T_{a})=\Delta_{CS}(T_{a}^{\prime})$ and $\Delta_{CS}(T_{b})=\Delta_{CS}(T_{b}^{\prime})$. This immediately leads to:
@@
    Thus, we have $\Delta_{CS}(T)=\Delta_{CS}(T^{\prime})$. However, by construction $T^{\prime}$ is either a Colless minimum or an echelon tree. Thus, we know that $T^{\prime}$ is $\Delta_{CS}$ minimal. This contradicts the assumption that $T$ is not $\Delta_{CS}$ minimal and thus completes the proof.
We conclude this section with the following corollary, which gives a recursive formula to count the number of $\Delta_{CS}$ minima.
@@
The proof is a direct consequence of Theorem 6 as all minimal trees are composed of two maximal pending subtrees which are also minimal, and these must have sizes that induce Colless or echelon partitions. In the even case, we need to avoid counting a combination of trees twice, which is why the simple sum is only considering the case $n_{a}>n_{b}$. In case $n_{a}=n_{b}=\frac{n}{2}$, on the other hand, every combination of two distinct trees is taken care of by the binomial coefficient, whereas every pair of equal trees is counted by $a\left(\frac{n}{2}\right)$, which explains the rightmost sum in the equation. This completes the proof. ∎
We conclude this section by noting that the sequence $(a_{n})_{n\geq 1}$ with $a_{n}=a(n)$, which starts with the values $\{1,1,1,1,2,2,1,1,3,6,6,4,4,3,1,1,4,13,25,26\}$ was not contained in the Online Encyclopedia of Integer Sequences OEIS [27], which indicates that it might not have appeared in other contexts before. It has been submitted to the OEIS in the course of this manuscript. Moreover, we want to mention that the reader can find more information on the set $\mathcal{M}(n)$ in [7], where all Colless partitions are characterized; so $\mathcal{M}(n)$ can simply be characterized by using those results and adding the echelon partition.
3.1.3 Further properties of $\Delta_{CS}$, $\Delta_{SC}$ and $N_{b}$